40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  Coating Ring-down Measurement Lab elog  Not logged in ELOG logo
Message ID: 311     Entry time: Fri Feb 17 13:44:47 2017
Author: Gabriele 
Type: Optics 
Category: Daily Progress 
Subject: Swapped picomotor and re-alignment 

Since I had recurrent problems with the picomotors used for QPD3, I swapped them with another Newport motorized mirror that was previously used in the Crackle1 experiment. This is the same model used for the other three QPD centering. Everything looks to be working fine now.

I also realigned all optical levers and swapped out an iris with a smaller one, to avoid beam clipping. All beam paths look clear now.

ELOG V3.1.3-